### v1

@@ 3.1 Preprocessing
The ABCD, UKB, and ABIDE datasets underwent initial preprocessing prior to analysis. In the ABCD study, a bandpass filter ranging from 0.009 to 0.08 Hz was applied to harmonize multi-site data [23]. In the UKB study, a bandpass filter ranging from 0.008 to 0.1 Hz was used to minimize the effects of heart rate variability and hemodynamic response delays [24]. In the ABIDE study, a finite impulse response (FIR) filter with a passband of 0.01–0.1 Hz [25] was employed to facilitate multi-site data integration. For the present study, additional preprocessing was performed on the resting-state BOLD data. The following preprocessing procedures were applied to each participant’s resting-state BOLD data: First, skull-stripping, slice-timing correction, and spatial normalization to MNI space were conducted using fMRIPrep ([26]). Subsequently, head motion artifacts and signals from white matter were removed using component-based noise correction (CompCor) [27]. The mean time-series signals from predefined regions of interest (ROIs) were extracted using the Nilearn package, with HCP-MMP1 symmetric ([28]) and Schaefer 400 ([29]) atlases employed for ROI definitions. Participants whose data contained voxels with NaN values, typically arising from the regression of unreliable white-matter signals, were excluded from the analysis. Additionally, to minimize spurious signal fluctuations due to scanner instability, the first 20 volumes of each time-series were discarded.
@@ 3.2 Dataset
Depression (ph9 score) 11.5$\pm$3.5 - -
Diagnosis (disorder/controls) - 141 (77/64) -
Fluid intelligence (z-score) - 0.01$\pm$0.91 -
@@ 3.3 Experimental Settings
For computational efficiency, pretraining was performed on a single A100 GPU, while training from scratch and fine-tuning were conducted using a single NVIDIA RTX 3090 GPU. The number of epochs was set to 100 for training from scratch. Efficient pretraining was achieved by determining the optimal number of epochs using weightwatcher ([31]), which identifies well-trained layers based on statistical learning theory. Optimal pretraining epochs were set to 1,000 for the HCP-MMP1 atlas and 400 for the Schaefer 400 atlas. Additional methodological details are provided in Appendix A.2.
3.4 Dividing frequencies
3.4.1 Frequency Dividing Strategy
@@
Let $w(f)$ be defined as:
@@
Then the spline multifractal function is:
@@
  • $\beta_{low}$ is the low-frequency scaling exponent.
@@
  • $\beta_{high}$ is the high-frequency scaling exponent.
  • $f_{2}$ is the knee frequency.
  • $s$ is the smoothness parameter.
To segment the resting-state fMRI data into distinct frequency bands, we applied both Lorentzian and spline multifractal equations to the power spectral density (PSD). These methods allowed us to identify knee frequencies ($f_{1}$, $f_{2}$) that demarcate the transitions between frequency bands and reveal their scale-free properties. The Lorentzian equation (Equation 2) was used to estimate $f_{1}$, the first knee frequency, by fitting the entire PSD. For frequencies above $f_{1}$, the spline multifractal equation (Equation 4) was applied to identify $f_{2}$, the second knee frequency. To ensure a smooth transition between different frequency regimes, we implemented cubic spline smoothing in the multifractal equation, preventing artifacts that could arise from abrupt transitions at the knee frequencies. The smoothing weight was adapted based on methodologies from ([32, 33]). The transition function $w(f)$ plays a crucial role in interpolating between two distinct power-law regimes. When $x\ll f_{2}$, $w(f)\approx 0$, ensuring that the function follows the lower-frequency power-law behavior characterized by $\beta_{\text{low}}$. Conversely, when $x\gg f_{2}$, $w(f)\approx 1$, shifting the function to the higher-frequency power-law behavior defined by $\beta_{\text{high}}$. In the intermediate range, $w(f)$ smoothly transitions between these regimes, with the parameter $s$ controlling the sharpness of the transition. This formulation enables a continuous and differentiable shift between scaling regimes while preserving the structural properties of the signal. These knee frequencies were used to segment the time-series data into three frequency bands: ultra-low frequency (below $f_{1}$), low frequency (between $f_{1}$ and $f_{2}$), and high frequency (above $f_{2}$). The segmentation process was implemented using the Finite impulse response (FIR) and Boxcar filter from the nitime package. The optimal filter types (FIR and Boxcar) were determined through systematic hyperparameter tuning on validation sets, evaluating their performance impact on the overall model. The Lorentzian function, having only two parameters, was efficiently fitted using the curve fit function from the SciPy package, which is well-suited for low-dimensional fitting. In contrast, the spline multifractal function, which involves five parameters, was fitted using the iminuit package, which is optimized for high-dimensional parameter estimation. In all three datasets, ANOVA analysis of $f_{1}$ and $f_{2}$ computed separately for each dataset, revealed a statistically significant difference between the two values ($p<0.001$), indicating that $f_{1}$ and $f_{2}$ reflect distinct frequency characteristics.
To assess the scale-free properties of these frequency bands, we computed $\beta$ values for each range, as shown in Equation 1. The results, summarized in Table 3, demonstrated significant differences in $\beta$ values across bands ($p<0.001$), supporting the hypothesis that each frequency component exhibits distinct scale-free dynamics. These results highlight that different frequency bands in the brain exhibit unique functional and topological properties. Specifically, lower frequencies are associated with long-range, global integration of brain activity, while higher frequencies correspond to more localized processing and rapid adaptation ([34, 35]). This finding reinforces the concept of multifractal behavior in brain signals, which has been previously observed in ECoG and EEG recordings ([20]). These $\beta$ values, denoted as $\beta_{ultralow}$, $\beta_{low}$, $\beta_{high}$, correspond to the ultra-low, low, and high-frequency bands, respectively.
3.4.2 Frequency Divison and Physiological Relevance
The values of $f_{1}$ and $f_{2}$ varied slightly across datasets (Table 2), yet consistently fell within specific ranges, reinforcing their physiological relevance. $f_{1}$ (0.046 Hz for ABCD, 0.062 Hz UKB datasets, and 0.018 Hz for ABIDE) corresponds to the time constant of an exponentially decaying function ([36]). Meanwhile, $f_{2}$ marks the transition to higher-frequency brain processes. $f_{2}$ was 0.078Hz for ABCD, 0.096Hz for UKB, and 0.030Hz for ABIDE. In ABCD and UKB, the $f_{2}$ values approaches the upper bound of conventional resting-state fMRI bandpass filtering (0.009-0.08Hz for ABCD, 0.008-0.1Hz for UKB). However, contemporary fMRI preprocessing pipelines commonly employ the Butterworth filter, which applies a gradual roll-off rather than an abrupt cutoff at upper bound (0.08Hz for ABCD, 0.1 Hz for UKB). This design minimizes phase distortion while effectively suppressing physiological noise from respiratory and cardiac ([37]). Consequently, the proximity of $f_{2}$ to upper bound of bandpass filtering does not imply a loss of meaningful higher-frequency information, as these signals are gradually attenuated rather than completely eliminated.
The significant differences in $\beta$ values across frequency bands further validate the frequency-dependent multifractal properties of fMRI signals. This demonstrates the potential of frequency-specific analyses to uncover multi-scale brain network organization and adaptive complexity in brain activity. Additionally, the reliability of $f_{1}$ and $f_{2}$ across different datasets emphasizes the generalizability of this approach, making it a powerful tool for further research into brain dynamics.
@@ 3.4.2 Frequency Divison and Physiological Relevance
The spline multifractal equation and the Lorentzian equation showed superior fitting performance across all datasets as depicted in Table 4. The consistent and high goodness-of-fit metrics across different datasets ($R^{2}>0.66$ for all cases) demonstrate the effectiveness of our dual-equation fitting approach, validating its reliability for characterizing the frequency-domain properties of resting-state fMRI data.
@@ 3.5.1 Model Architecture
Our model consists of two primary modules: the temporal module and the spatial module. These modules work synergistically to encode both the temporal and spatial characteristics of frequency-divided timeseries data, as shown in Figure 1 A. To ensure an effective balance between generalization and specificity, a parameter-sharing strategy is employed, as outlined below.
@@
As shown in Figure 1 B, the temporal and spatial modules operate on each frequency range, generating corresponding logits and attention matrices. To obtain the final prediction, the logits from the three frequency ranges are averaged and compared with the true label ($y_{true}$) using a task-specific loss function ($Loss_{task}$). In parallel, a spatial loss is computed to enforce distinct spatial connectivity patterns across frequency bands. The total loss function, which combines these two components, is defined as:
@@
In Equation 5, $Loss_{task}$ represents the cross-entropy loss for classification tasks or L1 loss for regression tasks. The $Loss_{spatial}$ term, as described in Section 9, emphasizes the distinct spatial connectivity patterns inherent to each frequency band. $\lambda$ was determined by hyperparameter tuning. The chosen $\lambda$ value and the filter type (described in Section 3.4.1) were shown in Table 5.
@@ 3.6.2 Using Original Form of BERT
In contrast, models like GPT are optimized for text generation tasks and do not emphasize inter-sequence contextual relationships, which are critical for our analysis. Unlike GPT’s causal masking—which is tailored to text generation and unidirectional context—BERT’s bidirectional attention allows it to capture long-range temporal dependencies from both past and future time points simultaneously, making it more suitable for modeling complex fMRI time series. Although some previous studies have used convolutional filters as a preprocessing step before entering signals into BERT ([40]), we chose to bypass this step. Although convolutional preprocessing can be computationally efficient, it may sacrifice the preservation of fine-grained temporal dynamics within the fMRI data, which could impact model performance (see Appendix D). By directly inputting the time-series data into the original BERT structure, we aimed to retain the full complexity of the temporal dynamics, allowing the model to effectively capture both short-term and long-term dependencies.
@@ 3.7.2 Spatial Loss
3.8 Pretraining Strategy Based on Communicability
@@ 3.8.1 Masking Strategy and Pretraining Loss
The goal of pretraining is to enable the model to learn the underlying structure of brain signal data in a self-supervised manner. To achieve this, we incorporated the concept of communicability into the pretraining process. A detailed definition of communicability used in this study is provided in Appendix A.1.
@@
Detailed model performance results (Appendix E.1) show that masking high-communicability nodes led to the most effective learning of the data structure, resulting in superior performance in downstream tasks.
@@
Pretraining involved two complementary masking strategies: temporal masking and spatial masking (Figure 4). Spatial masking involved nullifying the timeseries data of the top $k$ nodes with the highest communicability scores, where $k$ is less than the total number of ROIs. Temporal masking, on the other hand, initialized all ROI signals to zero within predefined time windows. Both strategies were applied to each frequency band (ultralow, low, and high), denoted as $\nu$.
@@ 3.9 Interpretability
Traditional attribution methods are not directly applicable to this model, as the relationship between attention matrices and target values is indirect. To address this challenge, we developed a novel approach inspired by GradCAM ([43]), specifically tailored to the spatial attention module described in Section 3.7.
@@
Statistically significant connections ($p<0.05$) and regions were analyzed, but the number of connections was too large to handle effectively. To address this, we selected only a few top connections based on the absolute values of Cohen’s d for further evaluation of their impact. An exhaustive list of all significant connections can be found in Appendix F.
@@ 3.10 Performance Comparison with Baseline Models
To evaluate the effectiveness of MBBN in learning spatiotemporal dynamics, we compared its performance with baseline models, including functional connectivity-based models (XGBoost [45], BNT [5]) and temporal dynamics models (BolT ([6]), vanilla BERT ([39])). These models were selected based on their relevance to brain dynamics interpretation, either through static functional connectivity or sequential processing of temporal data. While functional connectivity-based models primarily rely on static representations, temporal dynamics models capture sequential patterns but do not explicitly integrate spatial relationships. MBBN addresses this gap by combining both spatiotemporal characteristics using its frequency-dividing encoder and spatial attention module. Given that MBBN is an extension of vanilla BERT with additional architectural enhancements, a direct comparison between the two provides a robust evaluation of these modifications. While MBBN demonstrated superior performance across several downstream tasks, including the ABCD ADHD classification task, with significant improvements in accuracy and AUC scores, not all tasks showed a statistically significant advantage. Detailed performance comparison is described in Appendix E.1.
@@ 4 Discussion
Discussions should be brief and focused. In some disciplines use of Discussion or ‘Conclusion’ is interchangeable. It is not mandatory to use both. Some journals prefer a section ‘Results and Discussion’ followed by a section ‘Conclusion’. Please refer to Journal-level guidance for any specific requirements. This study introduces MBBN, a novel computational framework that integrates neuroscientific principles with advanced deep learning methodologies to analyze brain dynamics. MBBN employs a scale-free, frequency-based decomposition technique that encodes distinct frequency components as individual features, effectively capturing the multi-fractal dynamics of brain activity. Additionally, the model leverages a communicability-based masking strategy during pretraining, embedding the underlying network topology into its learned representations. Through a self-attention mechanism, MBBN combines biologically informed spatial relationships with temporal context, enabling the nuanced modeling of connectivity patterns and their evolution across multiple frequency ranges. This unified framework provides a robust platform for accurate, interpretable, and personalized predictions of cognitive and clinical outcomes.
Compared to previous transformer-based models for brain imaging data (e.g., [8, 7, 5, 6]), MBBN uniquely integrates the multi-fractal ([46, 47, 48]) and scale-free properties of neural signals ([36, 49]), essential for understanding the hierarchical and dynamic nature of brain activity ([50, 51, 52]). Unlike models relying solely on static or sequential representations ([5, 6]), or employing simplistic masking strategies ([7, 8]), MBBN’s biologically grounded approach—decomposing multifractal neural signals into interpretable frequency components and utilizing a communicability-driven masking scheme—enables it to effectively capture and reconstruct critical network hubs. By doing so, MBBN provides a more nuanced and biologically relevant representation of spatiotemporal connectivity.
@@
Crucially, correlation analysis confirms that the connectivity derived from MBBN’s self-attention mechanism represents a novel measure, capturing information distinct from and complementary to traditional functional connectivity (FC). To elucidate this, we directly compared attention-based connectivity with FC. This revealed a strong negative correlation when analyzing raw, non-frequency-divided signals (Spearman $\rho\approx-0.6$), indicating that attention-based connectivity captures fundamentally different, complementary information rather than merely replicating static FC patterns. Of note, the correlations became considerably weaker within specific frequency bands (ranging from $\rho\approx-0.1\backsim-0.2$) (Supplementary Table 2). These findings strongly suggest that MBBN successfully extracts unique, frequency-dependent connectivity patterns that are distinct from both traditional FC and the aggregated signal. This divergence not only validates our attention mechanism as providing a novel and meaningful measure of neural interaction but also underscores the importance of frequency-specific analysis for uncovering nuanced brain dynamics potentially missed by conventional methods. The biological plausibility of MBBN is further supported by its incorporation of the scale-free property, characterized by a power-law exponent ($\beta$), which differentiates brain states and regions based on their temporal correlation profiles ([2, 53, 54, 55, 56, 57]). Variations in $\beta$ values are associated with functional specialization and pathological alterations ([13, 14, 9, 58]), underscoring the importance of integrating these intrinsic properties into computational models. By incorporating communicability-based masking, MBBN systematically accounts for graph-structured brain data, addressing the limitations of random masking strategies ([59, 60]) and enhancing understanding of network-level interactions ([61, 62]).

### v2

@@ 3.1 Preprocessing
The ABCD, UKB, and ABIDE datasets underwent dataset-specific preprocessing to account for acquisition differences and multi-site harmonization requirements. In ABCD, a conservative bandpass filter (0.009-0.08 Hz) was applied following consortium protocols for adolescent populations ([20]). UKB employed a broader filter (0.008-0.1 Hz) optimized for adult populations to minimize heart rate variability effects ([21]). ABIDE used finite impulse response filtering (0.01-0.1 Hz) specifically designed for multi-site ASD data integration ([22]).
For the present study, additional preprocessing was performed on the resting-state BOLD data. The following preprocessing procedures were applied to each participant’s resting-state BOLD data: First, skull-stripping, slice-timing correction, and spatial normalization to MNI space were conducted using fMRIPrep ([23]). Subsequently, head motion artifacts and signals from white matter were removed using component-based noise correction (CompCor) ([24]). The mean time-series signals from predefined regions of interest (ROIs) were extracted using the Nilearn package, with HCP-MMP1 asymmetric ([25]) and Schaefer 400 ([26]) atlases employed for ROI definitions. Participants whose data contained voxels with NaN values, typically arising from the regression of unreliable white-matter signals, were excluded from the analysis. Additionally, to minimize spurious signal fluctuations due to scanner instability, the first 20 volumes of each time-series were discarded.
@@ 3.2 Dataset
Depression (PHQ-9 score) 11.5$\pm$3.5 - -
Depression (disorder/controls) - 3,113 (2,652/461) -
ADHD (disorder/controls) - 4,289 (2,439/1,850) -
ASD (disorder/controls) - - 141 (77/64)
@@ 3.3 Experimental Settings
For computational efficiency, pretraining was performed on a single A100 GPU, while training from scratch and fine-tuning were conducted using a single NVIDIA RTX 3090 GPU. The number of epochs was set to 100 for training from scratch. Efficient pretraining was achieved by determining the optimal number of epochs using weightwatcher ([28]), which identifies well-trained layers based on statistical learning theory. Optimal pretraining epochs were set to 1,000 for the HCP-MMP1 atlas and 400 for the Schaefer 400 atlas. Additional methodological details are provided in Supplementary Materials A.2.
3.4 Frequency-resolved decomposition based on scale-free principles
Brain activity exhibits multifractal properties with distinct scaling behaviors across frequency ranges, reflecting the hierarchical organization of neural networks from local circuits to global integration pathways. Conventional broadband connectivity analysis obscures these frequency-specific dynamics, potentially missing critical neurophysiological signatures of psychiatric disorders. To capture these frequency-dependent neural processes, we developed a systematic two-step fitting procedure to identify physiologically meaningful frequency boundaries (knee frequencies) that demarcate distinct scaling regimes in the power spectral density of resting-state fMRI signals. Following frequency decomposition using the identified knee frequencies, we confirmed that each of the three segments exhibited distinct power-law scaling behavior, as evidenced by significantly different $\beta$ values across frequency bands ($p<0.001$).
3.4.1 Step 1: Ultralow frequency boundary identification using Lorentzian fitting
We first applied the Lorentzian equation (Equation 2) to the entire power spectral density to identify $f_{1}$, the primary knee frequency that separates ultralow frequency activity from higher-frequency neural processes:
@@ 3.4.1 Step 1: Ultralow frequency boundary identification using Lorentzian fitting
where $A$ represents the amplitude parameter and $f_{1}$ denotes the characteristic frequency at which the power spectrum transitions from flat (low-frequency plateau) to steep decay (high-frequency roll-off). This Lorentzian model effectively captures the fundamental 1/f characteristics of neural signals while identifying the critical transition point. The fitting was performed using SciPy’s curve_fit function, which is well-suited for this two-parameter optimization problem. Physiologically, $f_{1}$ corresponds to the boundary between slow cortical potentials and faster network dynamics, typically associated with default mode network fluctuations and global integration processes.
3.4.2 Step 2: Low-high frequency boundary identification using spline multifractal modeling
For frequencies above $f_{1}$, we applied a sophisticated spline multifractal equation to identify $f_{2}$, the second knee frequency that distinguishes low-frequency network connectivity from high-frequency local processing:
To ensure smooth transitions between different scaling regimes and prevent fitting artifacts, we implemented a cubic spline weight function:
@@ 3.4.2 Step 2: Low-high frequency boundary identification using spline multifractal modeling
This weight function $w(f)$ provides a smooth, differentiable transition between two distinct power-law regimes. When $f\ll f_{2}$, $w(f)\approx 0$, ensuring the function follows low-frequency scaling behavior characterized by $\beta_{low}$. Conversely, when $f\gg f_{2}$, $w(f)\approx 1$, shifting to high-frequency scaling defined by $\beta_{high}$. The smoothness parameter $s$ controls the sharpness of this transition, preventing abrupt discontinuities that could introduce fitting artifacts.
The complete spline multifractal function then becomes:
@@
  • $\beta_{low}$ is the low-frequency scaling exponent (typically steeper, reflecting long-range correlations)
@@
  • $\beta_{high}$ is the high-frequency scaling exponent (typically flatter, reflecting local processing)
  • $f_{2}$ is the second knee frequency (transition point between network and local dynamics)
  • $s$ is the smoothness parameter controlling transition width
This five-parameter model was fitted using the iminuit package, specifically designed for high-dimensional parameter estimation with robust convergence properties. Physiologically, $f_{2}$ represents the transition from large-scale network connectivity to localized neural processing, typically corresponding to the boundary between task-negative and task-positive network activity.
3.4.3 Frequency band segmentation and filtering implementation
Based on the identified knee frequencies $f_{1}$ and $f_{2}$, we segmented each participant’s time-series data into three physiologically distinct frequency bands:
  1. 1.
    Ultra-low frequency band ($<f_{1}$): Captures slow cortical potentials, default mode network activity, and global integration processes
  2. 2.
    Low frequency band ($f_{1}$ to $f_{2}$): Represents intermediate-scale network connectivity and inter-regional communication
  3. 3.
    High frequency band ($>f_{2}$): Reflects local circuit dynamics, sensorimotor processing, and rapid cognitive operations
The segmentation was implemented using finite impulse response (FIR) filters from the nitime package for ultra-low and low-frequency bands, and Boxcar filters for high-frequency extraction. The choice of filter types was optimized through systematic hyperparameter tuning on validation sets, evaluating their impact on downstream classification performance. FIR filters were preferred for lower frequencies due to their superior phase preservation properties, while Boxcar filters provided computational efficiency for high-frequency components without compromising signal quality.
3.4.4 Validation of frequency-specific scaling properties
@@
To validate that our frequency decomposition captured distinct neurophysiological processes, we computed scaling exponents ($\beta$ values) for each frequency band using power-law fitting:
Statistical analysis revealed significant differences in $\beta$ values across all frequency bands ($p<0.001$), confirming that each band exhibits unique scale-free dynamics. Ultralow frequencies showed the flattest scaling (smallest $\beta$), consistent with long-range temporal correlations. High frequencies exhibited intermediate scaling, reflecting network-level integration. Low frequencies displayed the steepest scaling (largest $\beta$), characteristic of local, rapidly decorrelated processes. These findings support the hypothesis that frequency-specific decomposition captures fundamentally different aspects of brain organization, from global integration to local specialization.
@@ 3.5.1 Model Architecture
MBBN integrates temporal and spatial processing through a dual-module architecture that consists of the temporal module and the spatial module. These modules work synergistically to encode both the temporal and spatial characteristics of frequency-divided timeseries data, as shown in Figure 1 A. To ensure an effective balance between generalization and specificity, a parameter-sharing strategy is employed, as outlined below.
@@
As shown in Figure 1 B, the temporal and spatial modules operate on each frequency range, generating corresponding logits and attention matrices. To obtain the final prediction, the logits from the three frequency ranges are averaged and compared with the true label ($y_{true}$) using a task-specific loss function ($Loss_{task}$). In parallel, a spatial loss is computed to enforce distinct spatial connectivity patterns across frequency bands. The composite loss function enforces both task performance and frequency-specific spatial patterns:
@@
where task-specific loss ensures predictive accuracy while spatial loss encourages distinct connectivity patterns across frequency bands.
In Equation 5, $Loss_{task}$ represents the cross-entropy loss for classification tasks or L1 loss for regression tasks. The $Loss_{spatial}$ term, as described in Equation 9, emphasizes the distinct spatial connectivity patterns inherent to each frequency band. $\lambda$ was determined by hyperparameter tuning. The chosen $\lambda$ value and the filter type were shown in Table 6.
@@ 3.6.2 Using Original Form of BERT
In contrast, models like GPT are optimized for text generation tasks and do not emphasize inter-sequence contextual relationships, which are critical for our analysis. Unlike GPT’s causal masking—which is tailored to text generation and unidirectional context—BERT’s bidirectional attention allows it to capture long-range temporal dependencies from both past and future time points simultaneously, making it more suitable for modeling complex fMRI time series. Although some previous studies have used convolutional filters as a preprocessing step before entering signals into BERT ([31]), we chose to bypass this step. Although convolutional preprocessing can be computationally efficient, it may sacrifice the preservation of fine-grained temporal dynamics within the fMRI data, which could impact model performance (see Supplementary Materials D). By directly inputting the time-series data into the original BERT structure, we aimed to retain the full complexity of the temporal dynamics, allowing the model to effectively capture both short-term and long-term dependencies.
@@ 3.7.2 Spatial Loss
The negative logarithm amplifies subtle connectivity differences, ensuring the model learns distinct spatial patterns even when frequency bands show similar overall connectivity structures. We confirmed that this approach indeed derives neurobiologically plausible network properties (see Appendix B).
3.8 Communicability-based Pretraining Strategy
@@ 3.8.1 Masking Strategy and Pretraining Loss
The goal of pretraining is to enable the model to learn the underlying structure of brain signal data in a self-supervised manner. To achieve this, we incorporated the concept of communicability into the pretraining process. A detailed definition of communicability used in this study is provided in Supplementary Materials A.1.
High-communicability nodes act as critical hubs in brain networks, integrating information across distributed regions. By masking these influential nodes during pretraining, we force the model to learn their functional roles through contextual inference, similar to masked language modeling but tailored to brain network topology.
@@
Detailed model performance results (Supplementary Materials E.1) show that masking high-communicability nodes led to the most effective learning of the data structure, resulting in superior performance in downstream tasks.
@@
Pretraining involved two complementary masking strategies: temporal masking and spatial masking (Figure 4). Spatial masking involved masking the timeseries data of the top $k$ nodes with the highest communicability scores, where $k$ is less than the total number of ROIs. Temporal masking, on the other hand, initialized all ROI signals to zero within predefined time windows. Both strategies were applied to each frequency band (ultralow, low, and high), denoted as $\nu$.
@@ 3.9 Interpretability
Traditional attribution methods are not directly applicable to this model, as the relationship between attention matrices and target values is indirect. To address this challenge, we developed a GradCAM ([34])-inspired approach to identify frequency-specific connections driving psychiatric classification. This method highlights connections where gradient magnitude (importance) aligns with attention strength (activation), revealing clinically relevant biomarkers.
@@
Statistically significant connections ($p<0.05$) and regions were analyzed, but the number of connections was too large to handle effectively. To address this, we selected only a few top connections based on the absolute values of Cohen’s d for further evaluation of their impact. An exhaustive list of all significant connections can be found in Supplementary Materials F.
@@ 3.10 Performance Comparison with Baseline Models
To evaluate the effectiveness of MBBN in learning spatiotemporal dynamics, we compared its performance with baseline models, including functional connectivity-based models (XGBoost ([36]), BNT ([5])) and temporal dynamics models (BolT ([6]), vanilla BERT ([30])). These models were selected based on their relevance to brain dynamics interpretation, either through static functional connectivity or sequential processing of temporal data. While functional connectivity-based models primarily rely on static representations, temporal dynamics models capture sequential patterns but do not explicitly integrate spatial relationships. MBBN addresses this gap by combining both spatiotemporal characteristics using its frequency-dividing encoder and spatial attention module. Given that MBBN is an extension of vanilla BERT with additional architectural enhancements, a direct comparison between the two provides a robust evaluation of these modifications. While MBBN demonstrated superior performance across several downstream tasks, including the ABCD ADHD classification task, with significant improvements in accuracy and AUC scores, not all tasks showed a statistically significant advantage. Detailed performance comparison is described in Supplementary Materials E.1.
@@ 4 Discussion
This study presents MBBN, a novel deep learning framework that integrates biologically grounded principles with frequency-aware self-attention to model spatiotemporal brain dynamics. By leveraging scale-free and multifractal properties of neural signals, MBBN captures essential patterns of brain organization that are often overlooked by conventional transformer-based models. Additionally, MBBN employs communicability-based masking and frequency-aware self-attention to capture biologically meaningful connectivity patterns across multiple frequency ranges, enabling accurate and interpretable predictions of cognitive and clinical outcomes.
Unlike previous approaches that rely on broadband or static representations ([5, 6]), MBBN applies a frequency decomposition strategy to encode neural oscillations as discrete features, enabling a richer characterization of dynamic connectivity patterns. This frequency-resolved approach is particularly critical given that neural communication is mediated by distinct oscillatory dynamics across frequency bands, with high-frequency oscillations supporting local processing ([37]) and low-frequency rhythms facilitating long-range integration ([38]). Traditional broadband or narrowband analyses often obscure these nuanced patterns by averaging across spectrally diverse signals, potentially masking functionally specific effects. The MBBN framework overcomes these limitations by capturing such frequency-specific properties, allowing for a more precise and physiologically grounded understanding of brain network dynamics. Furthermore, our communicability-based ([39, 40]) masking method ensures that topological structures within brain networks are embedded into learned representations, improving both robustness and interpretability.
Importantly, MBBN-derived connectivity patterns complement rather than replicate traditional functional connectivity. While moderate anticorrelation existed at the broadband level, frequency-specific analyses revealed unique network dynamics (Supplementary Table 2), confirming that MBBN captures neurophysiologically distinct information critical for clinical interpretation.
@@
The biological plausibility of MBBN is further supported by its incorporation of the scale-free property, characterized by a power-law exponent ($\beta$), which differentiates brain states and regions based on their temporal correlation profiles ([2, 41, 42, 43, 44, 45]). Variations in $\beta$ values are associated with functional specialization and pathological alterations ([16, 17, 46, 47]), underscoring the importance of integrating these intrinsic properties into computational models. By incorporating communicability-based masking, MBBN systematically accounts for graph-structured brain data, addressing the limitations of random masking strategies ([48, 49]) and enhancing understanding of network-level interactions ([39, 50]).